RHIC Running: APEX running continued until 18:30 last night, thanks to the allowances from both experiments. Half the APEX program was completed. The remainder of yesterday's schedule will be added to next week's APEX.
Immediately after APEX, there was an access last night to allow PHENIX to remove their converter from the beampipe. During the access, Mike Brennan made adjustments to the stochastic cooling. The power supply for yo4 also needed to be replaced. STAR made use of the access as well.
A good physics store followed the access, starting at ~20:30. When filling the next store, beam losses dumped the beam, followed by a quench. While trying to recover from the quench, the QH8 magnet power supply in BtA failed at 4:55. .As of 9:00AM we are still trying to bring QH8 back on. While under LOTO for the QH8 work, the Polarimeter electronics at C15 was powered on. The RF station C amplifier was also replaced during this LOTO.
QH8 was brought back on, and the Landau cavity needed a new 5 watt amplifier. Colliding beams by noon. Radiation Interlock took the beam at 12:35. 90% BLAM trigger without any changes being made to the beam. Colliding beam came back by `2:00PM.
Scheduling Physicist: Mike Sivertz | Last updated: |
Email: sivertz@bnl.gov | |
Office Phone: 631-344-6102 | |
Digital Pager: 7528 or 4300 | |
Home Phone: 631-878-9020 |